# Noise-based reward-modulated learning

### v1

@@ 2.2.2 Noise-based learning rule
We can extend this by introducing noise into the neurons instead of the weights of the network, allowing localized perturbations within network layers. Given a layer transformation as in Equation (4), adding noise at the neuron level is represented as
$\tilde{x}^{l}=f\left(\tilde{h}^{l}+\xi^{l}\right)=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (13)
@@
where $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is the injected noise, $\beta$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$. We use the notation $\tilde{x}^{l}$ and $\tilde{h}^{l}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
@@
To formulate the learning rule for specific layer parameters $W^{l}\in W$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. To do so, we first rewrite the gradient term as
$\nabla p\left(W^{l}\right)=\nabla_{W^{l}}\log\pi_{\theta}\left(a\mid S_{t}\right)$ (14)
@@
and apply the chain rule of differentiation to break down the gradient with respect to $W^{l}$ as
$\nabla p\left(W^{l}\right)=\frac{\partial p}{\partial\tilde{h}^{l}}\frac{\partial\tilde{h}^{l}}{\partial W^{l}}=\nabla p\left(\tilde{h}^{l}\right)\tilde{x}^{l-1}\,.$ (15)
Building on the theory developed above, we derive the gradient $\nabla p(\tilde{h})$, leading to the layer-specific weight update rule
$W^{l}\leftarrow W^{l}+\eta\delta r_{t}\bar{\xi}^{l}{\rho}\tilde{x}^{l-1}$ (16)
with ${\rho}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$ and $\bar{\xi}^{l}=\ifrac{\xi^{l}}{||\xi^{l}||^{2}}$. Here, $\pi_{\tilde{W}}(a\mid S_{t})$ represents the network’s output when noise is injected into the neurons, while $\pi_{W}(a\mid S_{t})$ corresponds to the output of the noiseless network.
@@ 2.2.3 Learning from sparse rewards
In our derivation above, synaptic updates occur at every time step, implying that learning signals and rewards are received continuously. However, rewards are usually provided after completing a sequence of actions or upon reaching specific milestones, leading to more sparsely spaced feedback. To handle this scenario, we modify our learning rule to accommodate sparse rewards arriving at arbitrary times.
Starting from Equation (9), we redefine the synaptic weight update to include time explicitly
$W^{l}_{t+1}=W^{l}_{t}+\eta\;\delta r_{t}\;\bar{\xi}^{l}_{t}p_{t}\tilde{x}^{l-1}_{t}$ (17)
where $p_{t}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$. Next, we assume sparse rewards arriving at arbitrary times $\tau$, and thus synaptic updates only occur at this moment, denoted by $W^{l}_{\tau}$. The learning rule now accumulates information since the last reward was received, updating weights by integrating this information over time
$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}$ (18)
where $\tau_{0}$ represents the time when the previous reward was obtained.
We further represent this accumulated sum using an eligibility trace, which captures the history of neural activity and noise contributions since the last reward, defined as
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}\,.$ (19)
such that (18) may be expressed as
$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;e_{\tau}\,.$ (20)
Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (19), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (20), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights.
Eligibility traces act as a mechanism to connect past actions with future rewards [19]. They capture neural activity and other local variables over time, signaling potential synaptic changes. Upon receiving a reward, these traces are modulated by the reinforcement signal, resulting in synaptic updates. Some models view eligibility traces as decaying cumulative activity [21, 48]), while others treat them as a full activity history [32], which aligns with our formulation. The eligibility trace efficiently tracks neural information between rewards, facilitating the assignment of credit to past actions.
2.3 Neural network setup
At the core of the agent’s learning and decision-making process, we use a feedforward neural network with $L$ layers, where the transformation applied to each hidden layer is given by
$\tilde{x}^{l}$ $=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (21)
@@ 2.3 Neural network setup
$\tilde{y}$ $=s\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (22)
where $a^{l}$ and $x^{l}$ denote the layer pre-activation and output (with $x_{0}$ as the network input), $y$ is the network output, $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is noise added to the neurons, with $m^{l}$ representing the number of units in layer $l$. This is referred to as a “noisy pass”. The function $f(\cdot)$ represents a non-linear activation function applied to each layer. In our case, we use the LeakyReLU activation function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$, otherwise. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, converting the network’s output into action probabilities. We also define a “clean pass,” where no noise is present in the network, as
$x^{l}$ $=f\left(W^{l}x^{l-1}\right)$ (23)
$y$ $=s\left(W^{l}x^{l-1}\right)$ (24)
@@ 2.4 Experimental validation
We validate our approach across various environments and compare its performance with established baselines. Each environment presents an episodic problem, where the agent’s learning is guided solely by positive or negative rewards. We test in both instantaneous and delayed reward settings. In the delayed reward setting, a single reward is given after a sequence of actions, challenging the agent to assign credit retrospectively. For the instantaneous reward setting, we use the Reaching problem [18], while the delayed reward setting uses the Cartpole [1] and Acrobot [47] problems. We use implementations given by the libraries OpenAI Gym [3] and NeuroGym [33] for the different environments.
We compare NRL to two baselines: an exact-gradient version of NRL (similar to an actor-only variant of actor-critic which relies on a running average of past rewards as the prediction error), which serves as an ”optimal performance” benchmark and an RMHL approach. The first baseline uses BP for gradient computation, updating eligibility traces as
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\nabla_{W}\log\pi_{W}\left(a\mid S_{t}\right)\,$ (25)
@@
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\xi^{l}_{t}\tilde{x}^{l-1}_{t}\,.$ (26)
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau}$. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [37]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
@@ 3.1 Solving instantaneous and delayed reward problems
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section 3.2, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [18], visualized in Figure 1C, is an instantaneous reward problem that requires an agent to reach and stay at a target on a 1D ring by moving left, right, or staying still. At each step, the agent observes both the target’s position and its own, and it is rewarded based on its proximity to the target over fixed-duration trials. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1A and 1B, respectively.
@@
Second, the Acrobot problem [47], visualized in Figure 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the agent observes the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are based on the speed of completion, with a maximum time. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1D and 1E, respectively.
@@
The third and most challenging problem, the Cartpole problem [1], visualized in Figure 1I, is a delayed reward problem where an agent must balance a pole on a cart by moving left or right. At each time step, the agent observes the cart’s position and velocity, along with the pole’s angle and angular velocity, and acts accordingly. Performance is measured by the time the pole remains balanced, with a maximum time. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1G and 1H, respectively.
For all three tasks, we can observe that NRL reaches the performance level of BP whereas RMHL fails to do so.
3.2 Scalability to deeper networks
Figure 2: Performance on deeper networks – Acrobot. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Figure 3: Performance on deeper networks – Cartpole. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
@@ 3.2 Scalability to deeper networks
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers – a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Figure 2 displays the results for the Acrobot task and Figure 3 displays the results for the Cartpole tasks. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs. The results for neural networks with two hidden layers are shown in Figures 2A and B, and 3A and B, and those for networks with three hidden layers are presented in Figures 2C and D, and 3C and D.
@@
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases—an expected outcome due to the stochastic nature of the updates [20]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
@@ 3.3 Using only noisy passes
In NRL, ${p}_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, a completely noiseless network is not feasible in more biophysically realistic scenarios. Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero over time $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
Figure 4: Learning using only noisy passes – Acrobot. A: Clean pass approximation error. Each data point is computed with an absolute error, averaged over 500 timesteps using the Acrobot problem. B: Acrobot problem. Performance across trials, averaged over 5 runs, with mean, minimum, and maximum values displayed.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Figure 4A. Furthermore, Figure 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with the minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.

### v2

@@ 2.2.3 Learning with node level noise in neural network
Given a layer transformation as in Equation (3), adding noise at the neuron level is represented as
$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (9)
@@
where $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ is the injected noise at time $t$, $\sigma^{2}$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$. We use the notation $\tilde{x}^{l}_{t}$ and $\tilde{h}^{l}_{t}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
@@
To formulate the learning rule for specific layer parameters $W^{l}_{t}\in\{W^{1}_{t},\ldots,W^{L}_{t}\}$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$. To do so, we first rewrite the gradient term as
$\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ (10)
@@
and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right){{x}^{l-1\top}_{t}}.$ (11)
Notably, we may once again carry out the conversion from this gradient estimation step to a set of directional derivatives such that
[…]
where $i$ indexes over a set of repeated samples of the noise term, $\xi$.
Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (13)
[…]
@@ 2.2.4 Learning from sparse rewards
Our derivations so far assume that synaptic updates happen at every time step, implying a continuous stream of rewards and learning signals. In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones. To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
Let’s denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE, $\delta_{\tau_{m}}$, as
$\delta_{\tau_{m}}=r_{\tau_{m}}-\bar{r}_{\tau_{m}}.$ (14)
Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards.
$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right)$ (15)
While synaptic updates only occur at these specific reward times $\tau_{m}$, the learning rule continuously accumulates information between rewards. This information is integrated over time, allowing us to update weights based on what’s happened since the last reward. This leads to our modified learning rule
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\;e_{\tau_{m}}$ (16)
where
$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (17)
is an eligibility trace that acts as a mechanism to connect past actions with future rewards [23]. Eligibility traces capture neural activity and other local variables over time, signaling potential synaptic changes. Upon receiving a reward, these traces are modulated by the reinforcement signal, resulting in synaptic updates. Some models view eligibility traces as decaying cumulative activity [26, 54]), while others treat them as a full activity history [40], which aligns with our formulation. The eligibility trace efficiently tracks neural information between rewards, facilitating the assignment of credit to past actions.
Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (17), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (16), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights. This learning rule constitutes the core of our proposed noise-based reward-modulated Learning (NRL) and is used for all experiments presented in this paper.
2.3 Neural network architecture
The learning and decision-making capabilities of the system are modelled using a feedforward network with $L$ layers, representing interconnected populations of neurons. The transformation performed by each hidden layer at time $t$ can be expressed as
$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (18)
@@ 2.3 Neural network architecture
$\tilde{y}_{t}$ $=s\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (19)
where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer. This noisy propagation of activity is termed a “noisy pass.” The function $f(\cdot)$ is a non-linear activation function, which we implement as the LeakyReLU function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions. We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}x^{l-1\top}_{t}\right)$ and $y_{t}=s\left(W^{l}_{t}x^{l-1\top}_{t}\right)$.
@@ 2.4 Experimental validation
We validate our approach across a range of simulated environments, comparing it against established baseline learning methods. Each environment presents a problem with discrete episodes, where the system’s learning is guided solely by positive or negative rewards. We investigate both scenarios with immediate reward and those involving delayed reward. In the delayed reward scenarios, a single reward is provided after a sequence of actions, challenging our system to assign credit retrospectively to the actions that contributed to the outcome. For the instantaneous reward setting, we use the Reaching problem [22], while the delayed reward setting uses the Cartpole [1] and Acrobot [53] problems. We use implementations given by the libraries OpenAI Gym [5] and NeuroGym [41] for the different environments.
We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an ”optimal performance” benchmark, and an RMHL approach.
The first baseline, which we refer to as the exact-gradient method (BP baseline), is similar to an actor-only variant of actor-critic methods, relying on a running average of past rewards as the prediction error. For this baseline, we compute the exact policy gradients for the entire episode trajectory using backpropagation through time (BPTT) [57]. Given sparse rewards, the weights at each layer are updated at the end of each episode according to
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$ (20)
@@
$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{l}_{t}{\tilde{x}^{l-1\top}_{t}}\,.$ (21)
Similar to NRL’s noise-based learning rule defined in Equation (16), the synaptic weights for RMHL are then updated by modulating this eligibility trace with the reward prediction error.
In Table 1, we present the complete three learning rules (NRL, BPTT, and RMHL) in a single table for visualization and quick comparison. All updates are written in a per-layer form using $W^{l}$ and assume a delayed reward received at time $\tau_{m}$.
Table 1: Comparison of the learning rules.
Rule Update
NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,\tilde{x}^{\,l-1\top}_{t}$
BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$
RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,\tilde{x}^{\,l-1\top}_{t}$
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [44]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
@@ 3.1 Solving instantaneous and delayed reward problems
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section titled “Scalability to multi-layer networks”, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [22], visualized in Fig. 1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary. At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials. Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig 1A and Fig 1B, respectively.
@@
Second, the Acrobot problem [53], visualized in Fig 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are given based on the speed of completion, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1D and Fig 1E, respectively.
@@
The third and most challenging problem, the Cartpole problem [1], visualized in Fig 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right. At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly. Performance is measured by the time the pole remains balanced, with a maximum time allowed. Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1G and Fig 1H, respectively. For all three tasks, NRL reaches the performance level of BP, whereas RMHL fails to do so.
3.2 Scalability to multi-layer networks
@@
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Fig 2 displays the results for the Acrobot task and Fig 3 displays the results for the Cartpole tasks. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs. The results for neural networks with two hidden layers are shown in Fig 2A and Fig 2B, and 3A and B, and those for networks with three hidden layers are presented in Fig 2C and Fig 2D, and Fig 3C and Fig 3D.
Figure 2: Performance on deeper networks for the Acrobot. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Figure 3: Performance on deeper networks for the Cartpole. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
@@
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases – an expected outcome due to the stochastic nature of the updates [24]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
@@ 3.3 Using only noisy passes
In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, a completely noiseless network is not feasible in more biophysically realistic scenarios. Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig 4A. Furthermore, Fig 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
Figure 4: Learning using only noisy passes for the Acrobot. A: Clean pass approximation error. Each data point is computed with an absolute error, averaged over 500 timesteps using the Acrobot problem. B: Acrobot problem. Performance across trials, averaged over 5 runs, with mean, minimum, and maximum values displayed.